OpenStudy (anonymous):

Given Z = 5(cos98 + isin98) and W= 3(cos43 + isin43), find and simply Z/W. Round numerical entries in the answer to two decimal places.

OpenStudy (sagewilson):

This should help, ignore the red box.

OpenStudy (sagewilson):

Nevermind, I just realized that this doesn't go up to 98

OpenStudy (sagewilson):

I'll try to find another one.

OpenStudy (anonymous):

Thank you!

OpenStudy (sagewilson):

Okay you can use the calculator in google to fins the sine, cosine and tangent. The cos of 98 is -0.81928824529 and the sin of 98 is -0.57338187199 so Z is 5(-0.81928824529 + -0.57338187199)

OpenStudy (sagewilson):

W is 3(0.55511330152 + -0.83177474262) So it would look like this Z= -6.9633 or just -6.96 and W would be -0.829

OpenStudy (sagewilson):

Hope this helps.

zepdrix (zepdrix):

\[\large\rm Z=r_1(\cos a+i \sin a)\]\[\large\rm W=r_2(\cos b+i \sin b)\]Then,\[\large\rm Z/W=\frac{r_1}{r_2}(\cos[a-b]+i \sin[a-b])\]

zepdrix (zepdrix):

Divide the radial lengths, subtract the angles.

OpenStudy (jdoe0001):

and so... what did you get?